OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y =1/3x + 5 through (2, 1)

OpenStudy (anonymous):

A. y = 1/3x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

nvm

hero (hero):

1. Find the slope of the perpendicular line (The product of the slopes of lines perpendicular to each other is -1) 2. Insert the slope of the perpendicular line along with the point (2,-1) into y = mx + b 3. Solve for b: 4. Rewrite the point-slope formula of the perpendicular line, but include only the values of m and b

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

@Hero is A correct

OpenStudy (aripotta):

no

OpenStudy (anonymous):

@AriPotta then im screwed cause idk the answer

OpenStudy (aripotta):

perpendicular lines have negative reciprocal slopes. so if the one line has a slope of 1/3, what must be the slope of the perpendicular line?

OpenStudy (anonymous):

-1/3

OpenStudy (aripotta):

no. it would be -3

OpenStudy (anonymous):

ok whatever

OpenStudy (aripotta):

so the answer would be b. and it's not "whatever" if you're really trying to understand :/